Write a rule for the linear funtction whoes graph has slope m and contains the given point. m=3; (-2,2)
Answer by ankor@dixie-net.com(22740)

Write a rule for the linear function whoes graph has slope m and contains the given point. m=3; (-2,2)
:
Using the format: mx + b = y
:
replace m and y to find b
:
3(-2) + b = 2
-6 + b = 2
b = 2 + 6
b = 8
:
The equation: y = 3x + 8